# Maker-Breaker domination game critical graphs

### v1

@@ 5 MBD game critical graphs with small SMBD numbers
A graph $G$ is $2$-$\gamma_{\rm SMB}$-critical if and only if $G$ is the disjoint union of exactly one copy of $K_{1}$, at least one copy of $K_{1,n}$, $n\geq 2$ and possibly some copies of $K_{2}$.
Proof. Assume that $G$ is $2$-$\gamma_{\rm SMB}$-critical. Hence $\gamma_{\rm SMB}(G)=2$ and $\gamma_{\rm SMB}(G-e)=1$ for any edge $e\in E(G)$. Since $\gamma_{\rm SMB}(G)=2$ it follows from [3] that $G$ is exactly a graph that has either at least two strong support vertices or at least one strong support vertex and exactly one isolated vertex. Suppose first that $G$ has no isolated vertices. Since $\gamma_{\rm SMB}(G-e)=1$ for every $e\in E(G)$, it holds that $G-e$ has two isolated vertices for any edge $e\in E(G)$. Since $G$ has no isolated vertices, each component of $G$ is $K_{2}$. Hence $G=\ell K_{2}$ for some positive integer $\ell$, which leads to the contradiction as $2\neq\gamma_{\rm SMB}(\ell K_{2})=\infty$.
Now let $G$ be a graph with at least one isolated vertex and at least one strong support vertex. Since $\gamma_{\rm SMB}(G)\neq 1$, $G$ has exactly one isolated vertex. Since for any edge $e\in E(G)$ it holds that $G-e$ has at least two isolated vertices, every edge of $G$ is a pendant edge of $G$. Thus $G$ is the disjoint union of exactly one copy of $K_{1}$, at least one copy of $K_{1,n}$, $n\geq 2$ and possibly some copies of $K_{2}$.
Conversely, assume that $G$ is the disjoint union of exactly one copy of $K_{1}$, at least one copy of $K_{1,n}$, $n\geq 2$, and possibly some copies of $K_{2}$. Consider a D-game on $G$. Clearly, Dominator must first select the isolated vertex. Then Staller selects the centre of a $K_{1,n}$ (there is at least one such component), which allows her to win in her second move by choosing an unplayed leaf adjacent to the centre. Thus $\gamma_{\rm SMB}(G)=2$. Let $e$ be an edge of $G$. Clearly, $G-e$ has at least two isolated vertices. Therefore Staller can win the game in one move on $G-e$. This proves that $G$ is $2$-$\gamma_{\rm SMB}$-critical. $\square$
@@
A graph $G$ is $1$-$\gamma_{\rm SMB}^{\prime}$-critical if $\gamma_{\rm SMB}^{\prime}(G)=1$ and $\gamma_{\rm SMB}^{\prime}(G-e)<\gamma_{\rm SMB}^{\prime}(G)$ for any edge $e\in E(G)$. This is possible only when $G$ has at least one vertex and no edges. Thus we conclude that a graph $G$ is $1$-$\gamma_{\rm SMB}^{\prime}$-critical if and only if $G$ is a totally disconnected graph.
@@
A graph $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$ critical if and only if $G$ is the disjoint union of at least one copy of $K_{1,n},n\geq 2$, and possibly some copies of $K_{2}$.
Proof. Assume that $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$ critical. That is, $\gamma_{\rm SMB}^{\prime}(G)=2$ and $\gamma_{\rm SMB}^{\prime}(G-e)=1$ for any edge $e\in E(G)$. It is known from [6] that $\gamma_{\rm SMB}^{\prime}(G)=2$ if and only if $G$ has a strong support vertex. Also, $\gamma_{\rm SMB}^{\prime}(G-e)=1$ if and only if $G-e$ has an isolated vertex. This is possible only when every edge of $G$ is incident with a vertex of degree $1$. Thus we conclude that $G$ has at least one $K_{1,n},n\geq 2$, and possibly some copies of $K_{2}$.
Conversely, assume that $G$ is the disjoint union of at least one copy of $K_{1,n},n\geq 2$ and possibly some copies of $K_{2}$. Since $G$ has a strong support vertex, $\gamma_{\rm SMB}^{\prime}(G)=2$. It is clear that the removal of any edge from $G$ results in a graph having an isolated vertex. Thus $\gamma_{\rm SMB}^{\prime}(G-e)=1$ for any edge $e\in E(G)$. Thus $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$-critical. $\square$
6 Concluding remarks
The graph $G^{\prime}$ from the proof of Proposition 3.6 need not be connected. In fact, we believe that for any $k\geq 1$ there are no connected $k$-$\gamma_{\rm SMB}$-critical graphs. This is clearly true for $k=1$, as any 1-$\gamma_{\rm SMB}$-critical graph must contain at least two isolated vertices. Moreover, Theorem 5.1 implies that there do not exist connected $2$-$\gamma_{\rm SMB}$-critical graphs. For the general case we pose:
Conjecture 6.1
If $G$ is a connected graph, then $G$ is not $\gamma_{\rm SMB}$-critical.
In view of Proposition 4.4 and Theorem 4.5 we pose:
Open Problem 6.2
Find a characterization of $2$-$\gamma_{\rm MB}^{\prime}$-critical graphs.
@@ Acknowledgements
T. Dravec and S. Klavžar have been supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects N1-0285, N1-0218). A lot of work on this paper has been done during the Workshop on Games on Graphs II, June 2024, Rogla, Slovenia, the authors thank the Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia for supporting the workshop.

### v2

@@ 4 Critical graphs with small MBD numbers
A connected bipartite graph $G$ is $2$-$\gamma_{\rm MB}^{\prime}$-critical if and only if either $G=K_{2,m}$, $m\geq 3$, or $G=B_{n,m}$, $n,m\geq 3$.
A bit surprisingly, there exist also connected, non bipartite 2-$\gamma_{\rm MB}^{\prime}$-critical graphs. Let $G$ be a graph obtained from two disjoint copies of $K_{3}$ and a vertex $x$ by connecting $x$ with exactly one vertex in each copy of $K_{3}$. The graph $G$ is 2-$\gamma_{\rm MB}^{\prime}$-critical graphs. In view of this example and of Proposition 4.4 and Lemma 4.5, we conclude the section with:
Problem 4.8
Characterize $2$-$\gamma_{\rm MB}^{\prime}$-critical graphs.
5 SMBD game critical graphs
In this section, we introduce SMBD game critical graphs. By Lemma 2.3 (iii), (iv), the SMBD number of a graph never decreases by adding an edge to the graph. Hence, in this case we define critical graphs as follows.
Definition 5.1
If $G$ is a graph and $\tau\in\{\gamma_{\rm SMB},\gamma_{\rm SMB}^{\prime}\}$, then $G$ is $\tau$-critical, if $\tau(G)<\tau(G+e)$, for any edge $e\notin E(G)$.
Note that in view of Lemma 2.3 (iii) and (iv), if $G$ is $\tau$-critical, where $\tau\in\{\gamma_{\rm SMB},\gamma_{\rm SMB}^{\prime}\}$, then $\tau(G)<\infty$. If $G$ is $\tau$-critical and $\tau(G)=k$, where $\tau\in\{\gamma_{\rm SMB},\gamma_{\rm SMB}^{\prime}\}$, then we say that $G$ is a $k$-$\tau$-critical.
Let $\gamma_{\rm SMB}(G)=k$. Iteratively adding edges $e\notin E(G)$ with the property $\gamma_{\rm SMB}(G+e)=k$ we arrive at a $k$-$\gamma_{\rm SMB}$-critical graph. Hence, any graph $G$ with $\gamma_{\rm SMB}(G)=k$ is a spanning subgraph of a $k$-$\gamma_{\rm SMB}$-critical graph. Analogous conclusion holds for a $k$-$\gamma_{\rm SMB}^{\prime}$-critical graph. So such critical graphs clearly exists.
Proposition 5.2
If $G$ is a graph, then the following holds.
  1. (i)
    $G$ is $1$-$\gamma_{\rm SMB}$-critical if and only if $G=2K_{1}$ or $G=K_{n}\cup 2K_{1}$, $n\geq 1$.
  2. (ii)
    $G$ is $1$-$\gamma_{\rm SMB}^{\prime}$-critical if and only if $G=K_{n}\cup K_{1}$, $n\geq 1$.
Proof. (i) Assume that $G=2K_{1}$ or $G=K_{n}\cup 2K_{1}$, $n\geq 1$. Since $G$ has at least two isolated vertices, $\gamma_{\rm SMB}(G)=1$. Moreover, any edge $e\notin E(G)$ is incident to at least one of the isolated vertices. Therefore, $G+e$ contains at most one isolated vertex and hence $\gamma_{\rm SMB}(G+e)\geq 2$ for any $e\notin E(G)$. Thus $G$ is $1$-$\gamma_{\rm SMB}$-critical.
Conversely, assume that $G$ is $1$-$\gamma_{\rm SMB}$-critical, that is, $\gamma_{\rm SMB}(G)=1$ and $\gamma_{\rm SMB}(G+e)\geq 2$ for any $e\notin E(G)$. Hence $G$ has at least two isolated vertices, and for any $e\notin E(G)$, the graph $G+e$ has at most one isolated vertex. Therefore, if $e\notin E(G)$, then $e$ is incident to at least one of the isolated vertices. This is possible only when $G$ is either $2K_{1}$ or $K_{n}\cup 2K_{1}$ for $n\geq 1$.
(ii) Assume that $G$ is $K_{n}\cup K_{1}$ for $n\geq 1$. Since $G$ has an isolated vertex, $\gamma_{\rm SMB}^{\prime}(G)=1$. Moreover, any edge $e\notin E(G)$ is incident to the vertex of $K_{1}$. Therefore, $G+e$ has no isolated vertices and hence $\gamma_{\rm SMB}^{\prime}(G+e)\geq 2$ for any $e\notin E(G)$. Thus $G$ is $1$-$\gamma_{\rm SMB}^{\prime}$-critical.
@@ 5 SMBD game critical graphs
Conversely, assume that $G$ is $1$-$\gamma_{\rm SMB}^{\prime}$-critical, that is, $\gamma_{\rm SMB}^{\prime}(G)=1$ and $\gamma_{\rm SMB}^{\prime}(G+e)\geq 2$ for any $e\notin E(G)$. Thus $G$ has at least one isolated vertex and for any $e\notin E(G)$ it follows that $G+e$ has no isolated vertices. Therefore, $G$ has an isolated vertex $x$ such that any edge $e\notin E(G)$ is incident to $x$. This is possible only when $G=K_{n}\cup K_{1}$, $n\geq 1$. $\square$
To describe $2$-$\gamma_{\rm SMB}^{\prime}$-critical graphs and $2$-$\gamma_{\rm SMB}^{\prime}$-critical graphs, we define the following two families of graphs. A graph $G_{n}^{\prime}$ is a graph obtained from $K_{n}$ by attaching exactly two pendent vertices to a vertex of $K_{n}$, and a graph $G_{n}^{\prime\prime}$ is obtained from $K_{n}$ by respectively attaching exactly two pendent vertices to two different vertices of $K_{n}$. See Fig. 4.
Figure 4: Graphs $G_{5}^{\prime}\cup K_{1}$ and $G_{5}^{\prime\prime}$
For the proof of the next result we recall from [6, Proposition 4.4] that if $G$ is a graph with $\delta(G)\geq 1$, then $\gamma_{\rm SMB}(G)=2$ if and only if $G$ has at least two strong support vertices, and that $\gamma_{\rm SMB}^{\prime}(G)=2$ if and only if $G$ has a strong support vertex.
@@
If $G$ is a graph, then the following holds.
  1. (i)
    $G$ is $2$-$\gamma_{\rm SMB}$-critical if and only if $G=G_{n}^{\prime}\cup K_{1}$, $n\geq 1$, or $G=G_{n}^{\prime\prime}$, $n\geq 2$.
  2. (ii)
    $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$-critical if and only if $G=G_{n}^{\prime}$, $n\geq 1$.
Proof. (i) Assume that $G$ is $2$-$\gamma_{\rm SMB}$-critical. That is, $\gamma_{\rm SMB}(G)=2$ and $\gamma_{\rm SMB}(G+e)>2$ for any edge $e\notin E(G)$. Hence $G$ either has exactly one strong support vertex and exactly one isolated vertex, or $\delta(G)\geq 1$ and $G$ has exactly two strong support vertices.
First, assume that $G$ has exactly one strong support vertex $u$ with two leaf neighbors $u_{1},u_{2}$ and isolated vertex $v$. Since $\gamma_{\rm SMB}(G+e)>2$ for any edge $e\notin E(G)$, it follows that $G+e$ has either no strong support vertices or no isolated vertices. Hence any edge $e\notin E(G)$ is either incident with $v$ or with a vertex from $\{u_{1},u_{2}\}$. Thus $G=G^{\prime}\cup K_{1}$. Assume now that $G$ has exactly two strong support vertices and $\delta(G)\geq 1$. Since $G$ is critical, $G+e$ has at most one strong support vertex. This is possible only when any edge $e\notin E(G)$ is incident with at least one pendant vertex of $G$ that eliminates at least one strong support vertex. Thus $G=G_{n}^{\prime\prime}$.
Conversely assume that $G=G_{n}^{\prime}\cup K_{1}$, $n\geq 1$, or $G=G_{n}^{\prime\prime}$, $n\geq 2$. Then $G$ has either a strong support vertex and an isolated vertex, or two strong support vertices. Therefore $\gamma_{\rm SMB}(G)=2$. Since any edge $e\notin E(G)$ is incident to either the isolated vertex or a pendant vertex, the graph $G+e$ has either at most one strong support vertex and no isolated vertices, or one isolated vertex and no strong support vertices. Therefore $\gamma_{\rm SMB}(G+e)>2$. Thus $G$ is $2$-$\gamma_{\rm SMB}$-critical.
(ii) Assume that $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$-critical, that is, $\gamma_{\rm SMB}^{\prime}(G)=2$ and $\gamma_{\rm SMB}^{\prime}(G+e)>2$ for any edge $e\notin E(G)$. Hence $\delta(G)=1$ and $G$ has exactly one strong support vertex. Moreover, any edge $e\notin E(G)$ is incident to a pendant vertex so that $G+e$ has no strong support vertices. Hence $G=G_{n}^{\prime}$.
Conversely assume that $G=G_{n}^{\prime}$. Since $G$ has a strong support vertex with exactly two leaf neighbors, $\gamma_{\rm SMB}^{\prime}(G)=2$. As any edge $e\notin E(G)$ is incident with at least one of the pendant vertices, $G+e$ has no strong support vertices and hence $\gamma_{\rm SMB}^{\prime}(G+e)>2$. Thus $G$ is $2$-$\gamma_{\rm SMB}^{\prime}$-critical. $\square$
Recall that $G|X$ denotes the graph $G$ in which vertices from $X\subseteq V(G)$ are considered as being already dominated. For $v\in V(G)$, let $G_{v}$ be a graph obtained from $G|N[v]$ by deleting all dominated vertices $x\in V(G|N[v])$ with $N[x]\subseteq N[v]$. Assume now that $G$ is 2-$\gamma_{\rm SMB}$-critical, so that in view of Theorem 5.3, $G$ is $K_{1}\cup G_{n}^{\prime}$ for some $n\geq 1$, or $G_{n}^{\prime\prime}$ for $n\geq 2$. Then there exists a vertex $v\in V(G)$ (an isolated vertex or a strong support vertex of $G$) such that $G_{v}$ is 2-$\gamma_{\rm SMB}^{\prime}$-critical. We wonder if this holds also for $k$-$\gamma_{\rm SMB}$-critical graphs with $k\geq 3$:
Question 5.4
Let $G$ be a $k$-$\gamma_{\rm SMB}$-critical graph, $k\geq 3$. Is it true that there exists a vertex $v\in V(G)$ such that $G_{v}$ is $\gamma_{\rm SMB}^{\prime}$-critical?
@@ Acknowledgements
T. Dravec and S. Klavžar have been supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects N1-0285, N1-0355). A lot of work on this paper has been done during the Workshop on Games on Graphs II, June 2024, Rogla, Slovenia, the authors thank the Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia for supporting the workshop.